roberta has 4 dollars in dimes and quarters. She has 5 more dimes than quarters.In total, roberta has 25 coins

Answer by CPhill(1959)

Here's how we can solve this problem:
Let $q$ be the number of quarters and $d$ be the number of dimes.
Since Roberta has 5 more dimes than quarters, we have the equation:
$d = q + 5$
The total number of coins is 25, so we have:
$q + d = 25$
Substituting the first equation into the second equation, we get:
$q + (q + 5) = 25$
Combining like terms, we have:
$2q + 5 = 25$
Subtracting 5 from both sides, we get:
$2q = 20$
Dividing both sides by 2, we find:
$q = 10$
Substituting this value back into the equation $d = q + 5$, we find:
$d = 10 + 5 = 15$
So, Roberta has 10 quarters and 15 dimes.
We can check our answer by verifying that the total amount of money is $4:
$0.25(10) + $0.10(15) = $2.50 + $1.50 = $4.00
Therefore, our answer is correct.